A package of 12 recordable CDs contains 3 which are defective. If 4 CDs are selected at random, find the probability that:
a. No defective CDs are found.
b. 1 of the 4 CDs is defective.
c. 3 of the 4 CDs are defective.
Not a text book question. Thank you for your help :)

Answer by stanbon(57380)

A package of 12 recordable CDs contains 3 which are defective. If 4 CDs are selected at random, find the probability that:
a. No defective CDs are found.
b. 1 of the 4 CDs is defective.
c. 3 of the 4 CDs are defective.
------------
P(defective)= 3/12 = 1/4; P(non defective)=3/4
----------------------
a. P(No defective CDs are found)= (3/4)^4
b. P(1 of the 4 CDs is defective)= 4C1(1/4)(3/4)^3 = 0.421875...
c. P(3 of the 4 CDs are defective)=4C3(1/4)^3(3/4)= 0.421875...
--------------
Cheers,
Stan H.